OpenStudy (anonymous):

this is a real question, can you teach me how to do it if you want?x • x • x • x • x = ?

OpenStudy (saifoo.khan):

\[\Large x^{1+1+1+1+1} = x^5\]

OpenStudy (turingtest):

...because the rule is\[\large x^a\cdot x^b=x^{a+b}\]and\[\large x=x^1\]so between these two we have\[\large x\cdot x\cdot x\cdot x\cdot x=x^1\cdot x^1\cdot x^1\cdot x^1\cdot x^1\cdot x^1=x^{1+1+1+1+1}=x^5\]

OpenStudy (turingtest):

*there are only supposed to be 5 x^1's there in the last line...